Suppose NBA player weights (in pounds) are N(221, 15^2 ).

(a) Find the weight such that 20% of players weigh less than that weight.

(Can you please explain in detail how to use the Z-Table with this one?)

(b) A random group of 5 NBA players (still with weights from N(221, 152 )) cross a playground bridge together, even though its breaking strength is only 1000 pounds. What is the probability that it breaks?
Answer #1

Ans:

a)

P(Z<=z)=0.20

z=normsinv(0.20)

z=-0.84

(see the row and column corresponding to a cumulative probability of 0.20,it will be a negative z value with row -0.8 and column 0.04.

Now,

x=221-0.84*15=208.4

b)

sampling distribution of sample sum:

mean=221*5=1105

standard deviation=sqrt(5)*15=33.541

z=(1000-1105)/33.541

z=-3.13

P(breaks)=P(z>-3.13)=0.9991